# HaluMem: Evaluating Hallucinations in Memory Systems of Agents

### v2

@@ 2.2 Evaluation Hallucinations in Memory Systems
Table 2: Comparison of Hallucination Evaluation Benchmarks
@@
Evaluation Tasks
Memory Extraction,
Memory Updating,
Memory QA
Multiple Choice
QA,
Summarization,
Generation
QA,
Memory Recall
Generation,
Classification
Memory Type
Persona,
Event,
Relationship
Persona
Persona,
Event
Persona,
Event
Persona
@@
In contrast, research on memory hallucinations remains at an early stage. Representative studies include LoCoMo, LongMemEval, PrefEval, and PersonaMem (Table 2). The early benchmark LoCoMo [maharana2024evaluating] focuses on memory retention under long-context settings. It adopts an end-to-end evaluation paradigm, assessing models through question answering, summarization, and generation tasks to test factual recall and event tracking over ultra-long texts. Although the dataset is relatively large (about 7.5k questions), it only covers a time span of several months and lacks an explicit memory updating mechanism, primarily reflecting the model’s capability for static information retention. Subsequently, LongMemEval [wulongmemeval] extends this framework by introducing metrics such as Information Retention Rate and Memory Recall Accuracy, covering approximately 2.5 years of multi-turn dialogue and incorporating explicit memory updates to quantify knowledge consistency across time. This represents a shift from static evaluation toward dynamic memory modeling. In the direction of personalization, PrefEval [zhao2025do] evaluates a model’s ability to maintain and follow user preferences over long-term interactions, using generation and classification tasks to assess preference consistency, though it remains limited to persona-level memory. PersonaMem [jiang2025know] further constructs simulated user personas and event histories, employing multiple-choice evaluations to assess persona consistency, traceability, and update accuracy. With a longer time span (on the order of years), it provides a more representative benchmark for personalized long-term memory assessment.
Although these studies provide preliminary frameworks for evaluating memory-related capabilities, they remain predominantly end-to-end and fail to localize hallucinations at the operational level within memory systems.
@@ 3 Problem Definition
Figure 2: Comparison between HaluMem and existing hallucination evaluation methods for memory systems.
@@ 5 Evaluation Framework of HaluMem
For each user, the session-level evaluation procedure of HaluMem is defined as follows: (1) Dialogue sessions $D^{1},D^{2},\dots,D^{S}$ are sequentially fed, in chronological order, into the memory system $S$. (2) If the current session $D^{s}$ contains reference memory points or QA tasks [^6] The evaluation of QA tasks follows the conventional end-to-end procedure. Therefore, in Figure 4, we only illustrate the simplified testing process for hallucination evaluation in memory extraction and memory updating., the corresponding evaluation process (extraction, updating, or question answering) is triggered immediately after $S$ completes processing that session, and the results are recorded. (3) After processing all sessions, the metrics of the three categories of tasks are aggregated to obtain the overall system performance.
Figure 4: Hallucination evaluation process
@@ 5.1 Memory Extraction
where $N_{\mathrm{extract}}=|\widehat{M}^{\mathrm{ext}}_{s}|$, and $M_{T}$ denotes the set of target memories that match the reference ones.
@@ 5.2 Memory Updating
The memory updating task evaluates whether the system can correctly modify, merge, or replace existing memories during new dialogues so that consistency is maintained without introducing hallucinations. For each dialogue session $D^{s}$ that contains annotated updates, the gold update set is defined as
$G^{\mathrm{upd}}_{s}=\{(m^{\mathrm{old}}\rightarrow m^{\mathrm{new}})\},$
The system output is denoted as $\widehat{G}^{\mathrm{upd}}_{s}$.
@@
Memory Update Accuracy $=\frac{N_{\mathrm{correct\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$ (4)
Memory Update Hallucination Rate $=\frac{N_{\mathrm{wrong\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$
Memory Update Omission Rate $=\frac{N_{\mathrm{missed\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$
@@ 5.3 Memory Question Answering
$\mathrm{QA\text{-}Accuracy}$ $=\frac{N_{\mathrm{correct}}}{N_{\mathrm{total}}},$ (5)
$\mathrm{QA\text{-}Hallucination}$ $=\frac{N_{\mathrm{hallucinated}}}{N_{\mathrm{total}}},$
$\mathrm{QA\text{-}Omission}$ $=\frac{N_{\mathrm{omitted}}}{N_{\mathrm{total}}},$
where $N_{\mathrm{total}}$ denotes the total number of test questions, $N_{\mathrm{correct}}$ represents the number of questions that are correctly answered, $N_{\mathrm{hallucinated}}$ indicates the number of questions that contain fabricated or incorrect information, and $N_{\mathrm{omitted}}$ refers to the number of questions that are left unanswered due to missing memories.
@@ 6.1 Experimental Setup
We conducted a comprehensive evaluation of several state-of-the-art memory systems on HaluMem, including Mem0 (both standard and graph versions) [chhikara2025mem0], Memobase [memobase], Supermemory [supermemory], and Zep [rasmussen2025zep]. [^7] MemOS will also be included in the evaluation as soon as possible. Each memory system was independently evaluated in two subsets, HaluMem-Medium and HaluMem-Long, with efforts made to ensure consistent parameter configurations across evaluations.
@@ 6.2.1 Overall Evaluation on HaluMem
Table 3: Evaluation results of all memory systems on HaluMem. “R” denotes Recall, “Target P” denotes Target Memory Precision, “Acc.” denotes Accuracy, “FMR” denotes False Memory Resistance, “C” denotes Correct Rate (Accuracy), “H” denotes Hallucination Rate, and “O” denotes Omission Rate. The values in parentheses in the “Target P” and “Acc.” columns represent the number of extracted memories. Color scale reflects performance (red = worse, green = better); Best values in bold.
Dataset System Memory Integrity Memory Accuracy Memory Updating Question Answering
@@
Medium Mem0 42.91% 65.03% 86.26%(10556) 60.86%(16291) 56.80% 25.50% 0.45% 74.02% 53.02% 19.17% 27.81%
Mem0-Graph 43.28% 65.52% 87.20%(10567) 61.86%(16230) 55.70% 24.50% 0.26% 75.24% 54.66% 19.28% 26.06%
Memobase 14.55% 25.88% 92.24%(5443) 32.29%(17081) 80.78% 5.20% 0.55% 94.25% 35.33% 29.97% 34.71%
Supermemory 41.53% 64.76% 90.32%(14134) 60.83%(22551) 51.77% 16.37% 1.15% 82.47% 54.07% 22.24% 23.69%
Zep - - - - - 47.28% 0.42% 52.31% 55.47% 21.92% 22.62%
Long Mem0 3.23% 11.89% 88.01%(1134) 46.01%(2433) 87.65% 1.45% 0.03% 98.51% 28.11% 17.29% 54.60%
Mem0-Graph 2.24% 10.76% 87.32%(785) 41.26%(1866) 88.36% 1.47% 0.04% 98.40% 32.44% 21.82% 45.74%
Memobase 6.18% 14.68% 88.56%(3077) 25.61%(11795) 85.39% 4.10% 0.36% 95.38% 33.60% 29.46% 36.96%
Supermemory 53.02% 70.73% 85.82%(24483) 29.71%(77134) 36.86% 17.01% 0.58% 82.42% 53.77% 22.21% 24.02%
Zep - - - - - 37.35% 0.48% 62.14% 50.19% 22.51% 27.30%
  • Note: since Zep does not provide a Get Dialogue Memory API, it is not possible to calculate Memory Integrity or Memory Accuracy. For details, see Appendix B.
@@
Overall, almost all memory systems perform worse on HaluMem-Long than on HaluMem-Medium, with Mem0, Mem0-Graph, and Memobase showing particularly notable declines. Notably, except for Supermemory and Zep, all other systems extract significantly fewer memories on HaluMem-Long than on HaluMem-Medium. This indicates that future memory systems need to improve their ability to process irrelevant information and distinguish between high- and low-value memories.
In the memory extraction task, all systems achieve recall (R) rates below 60% in terms of memory integrity, indicating that many reference memory points are not extracted. The better performance in weighted memory recall (weighted R) compared to recall suggests that the systems can prioritize important memory points. Regarding memory accuracy, all systems have accuracy (Acc.) below 62%, reflecting a high proportion of hallucinations, although performance on Target P is relatively good. Supermemory performs the worst on FMR because it tends to extract excessive information without effectively filtering distractions or unhelpful content. Other systems adopt more conservative strategies and thus perform better in FMR. In summary, future memory systems should strike a balance among coverage of important memories, extraction accuracy, and resistance to interference, aiming for both high quality and reliability in memory retrieval.
In the memory updating task, all systems achieve correct update rates below 50%, and except for Supermemory, their performance drops considerably on HaluMem-Long. Systems showing better performance in memory integrity also tend to exhibit higher update accuracy, but all suffer omission rates above 50%. This issue primarily stems from insufficient coverage in memory extraction: when the pre-update memories are not extracted, related updates cannot be properly processed. Moreover, the fact that all systems exhibit hallucination rates below 2% does not necessarily imply strong hallucination suppression, since very few samples actually enter the update stage. Overall, current systems face a clear bottleneck in memory updating: the extraction and updating stages lack stable linkage, resulting in low accuracy and high omission rates.
In the memory question-answering task, the best-performing systems are also those that perform well in memory integrity and memory updating, further highlighting the crucial role of memory extraction. For example, Mem0 and Mem0-Graph show clear performance declines on HaluMem-Long compared to HaluMem-Medium, which strongly correlates with their substantial reduction in extracted memory points. However, all systems achieve answer accuracies below 56%, with both hallucination rate and omission rate remaining high, and their overall performance further decreases on HaluMem-Long. This demonstrates that current memory systems’ QA performance depends heavily on the sufficiency and accuracy of upstream memory extraction, and remains prone to factual deviation and memory confusion under interference or extended context conditions.
6.2.2 Performance on Different Memory Types
@@
Table 4 reports the extraction accuracy of each memory system for event, persona, and relationship memories, which include all memory points from both the memory extraction and updating tasks while excluding distractor memories. Zep achieve the best performance on HaluMem-Medium, whereas Supermemory performs best on HaluMem-Long. However, Mem0, Mem0-Graph, and Memobase show a marked decline in long-context scenarios, suggesting difficulty in consistently capturing valuable information in complex dialogues. Supermemory is the only one whose performance on HaluMem-Long surpasses that on HaluMem-Medium, probably because it extracts a larger number of memory points in the long-context condition. Across memory types, Persona memories yield slightly higher accuracy, indicating that static personal traits are easier to capture, whereas understanding event dynamics and relationship changes remains challenging. Overall, all systems still show low performance across the three memory categories, indicating significant limitations in current memory modeling.
@@ 6.2.3 Performance on Different Question Types
Figure 5 illustrates the performance of different memory systems across six categories of questions. Overall, the accuracy of all memory systems remains relatively low across most categories, indicating substantial room for improvement. The Mem0 series and Memobase show significantly poorer performance on HaluMem-Long compared to HaluMem-Medium, suggesting a notable degradation under ultra-long context conditions. In contrast, SuperMemory and Zep demonstrate relatively stable behavior and achieve consistently superior overall performance on both datasets. Furthermore, all memory systems perform comparatively better on memory boundary and memory conflict questions, indicating their capability to effectively recognize unknown or misleading information and respond correctly. However, their performance deteriorates substantially on multi-hop inference, dynamic update, and generalization & application questions, suggesting that current memory systems still struggle with complex reasoning and preference tracking.
@@ 6.2.4 Efficiency Analysis of Memory Systems
Table 5 shows the time consumption of all memory systems during the evaluation process for dialogue addition and memory retrieval, as well as their total runtime. Overall, dialogue addition requires substantially more time than memory retrieval, indicating that the write stage is the primary computational bottleneck. Enhancing the efficiency of memory extraction and updating is thus crucial for improving interactive performance. On HaluMem-Medium, Supermemory performs best in both dialogue addition and total runtime, while Mem0 shows the best retrieval efficiency. However, the dialogue addition time of Mem0 and Mem0-Graph exceeds 2700 minutes, revealing their low processing efficiency during dialogue ingestion and memory construction. On HaluMem-Long, the dialogue addition time for Mem0, Mem0-Graph, and Memobase decreases, mainly because the number of processed memory points is reduced rather than due to performance improvement. In contrast, Supermemory extracts a substantially larger number of memory points, resulting in the highest time cost.

### v3

@@ 2.2 Evaluation Hallucinations in Memory Systems
Table 2: HaluMem vs. Existing End-to-End Benchmarks for Memory System Evaluation
@@
Evaluation Tasks Memory Extraction, Memory Updating, Memory QA Multiple Choice QA, Summarization, Generation QA, Memory Recall Generation, Classification
Memory Type Persona, Event, Relationship Persona Persona, Event Persona, Event Persona
@@
In contrast, research on hallucinations in memory systems is still in its infancy. Existing benchmarks such as LoCoMo, LongMemEval, PrefEval, and PersonaMem (Table 2) focus on overall memory system performance rather than hallucination-specific evaluation. The early benchmark LoCoMo [maharana2024evaluating] focuses on memory retention under long-context settings. It adopts an end-to-end evaluation paradigm, assessing models through question answering, summarization, and generation tasks to test factual recall and event tracking over ultra-long texts. Although the dataset is relatively large (about 7.5k questions), it only covers a time span of several months and lacks an explicit memory updating mechanism, primarily reflecting the model’s capability for static information retention. Subsequently, LongMemEval [wulongmemeval] extends this framework by introducing metrics such as Information Retention Rate and Memory Recall Accuracy, covering approximately 2.5 years of multi-turn dialogue and incorporating explicit memory updates to quantify knowledge consistency across time. This represents a shift from static evaluation toward dynamic memory modeling. In the direction of personalization, PrefEval [zhao2025do] evaluates a model’s ability to maintain and follow user preferences over long-term interactions, using generation and classification tasks to assess preference consistency, though it remains limited to persona-level memory. PersonaMem [jiang2025know] further constructs simulated user personas and event histories, employing multiple-choice evaluations to assess persona consistency, traceability, and update accuracy. With a longer time span (on the order of years), it provides a more representative benchmark for personalized long-term memory assessment.
Existing benchmarks primarily adopt holistic, end-to-end evaluations that treat memory systems as black boxes, making hallucinations only indirectly observable through final task performance rather than explicitly attributable to specific memory operations. In contrast, HaluMem is the first benchmark specifically designed to evaluate hallucinations in memory systems, enabling fine-grained, operation-level analysis and filling a critical gap left by prior memory evaluation benchmarks.
@@ 3 Problem Definition
Figure 2: Comparison between HaluMem and existing benchmarks for memory systems.
@@ 5 Evaluation Framework of HaluMem
For each user, the session-level evaluation procedure of HaluMem is defined as follows: (1) Dialogue sessions $D^{1},D^{2},\dots,D^{S}$ are sequentially fed, in chronological order, into the memory system $S$. (2) If the current session $D^{s}$ contains reference memory points or QA tasks, the corresponding evaluation process (extraction, updating, or question answering) is triggered immediately after $S$ completes processing that session, and the results are recorded. (3) After processing all sessions, the metrics of the three categories of tasks are aggregated to obtain the overall system performance.
Figure 4: Hallucination evaluation process.
@@ 5.1 Memory Extraction
where $N_{\mathrm{extract}}=|\widehat{M}^{\mathrm{ext}}_{s}|$, and $M_{T}\subset\widehat{M}^{\mathrm{ext}}_{s}$ denotes the set of target memories that match the reference ones.
@@
Memory Extraction F1
We additionally report an F1 score to measure the overall performance of the memory extraction task by jointly considering completeness and correctness. Memory Recall ($R_{\text{mem}}$) is used as the recall term, while Target Memory Precision ($P_{\text{tgt}}$) is used as the precision term. The F1 score is defined as:
$\mathrm{F1}_{\text{mem}}=\frac{2\,R_{\text{mem}}\,P_{\text{tgt}}}{R_{\text{mem}}+P_{\text{tgt}}}$ (4)
@@ 5.2 Memory Updating
The memory updating task evaluates whether the system can correctly modify, merge, or replace existing memories during new dialogues so that consistency is maintained without introducing hallucinations. For each dialogue session $D^{s}$ that contains annotated updates, the gold update set is defined as $G^{\mathrm{upd}}_{s}=\{(m^{\mathrm{old}}\rightarrow m^{\mathrm{new}})\}$. The system output is denoted as $\widehat{G}^{\mathrm{upd}}_{s}$.
@@
Memory Updating Accuracy $=\frac{N_{\mathrm{correct\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$ (5)
Memory Updating Hallucination Rate $=\frac{N_{\mathrm{wrong\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$
Memory Updating Omission Rate $=\frac{N_{\mathrm{missed\text{-}upd}}}{N_{\mathrm{target\text{-}upd}}},$
@@ 5.3 Memory Question Answering
Memory QA Accuracy $=\frac{N_{\mathrm{correct\text{-}qa}}}{N_{\mathrm{total\text{-}qa}}},$ (6)
Memory QA Hallucination Rate $=\frac{N_{\mathrm{wrong\text{-}qa}}}{N_{\mathrm{total\text{-}qa}}},$
Memory QA Omission Rate $=\frac{N_{\mathrm{missed\text{-}qa}}}{N_{\mathrm{total\text{-}qa}}},$
where $N_{\mathrm{total\text{-}qa}}$ denotes the total number of questions, $N_{\mathrm{correct\text{-}qa}}$ denotes the number of correctly answered questions, $N_{\mathrm{wrong\text{-}qa}}$ denotes the number of questions answered with fabricated or incorrect information, and $N_{\mathrm{missed\text{-}qa}}$ refers to the number of questions that are left unanswered due to missing memories.
@@ 6.1 Experimental Setup
We conducted a comprehensive evaluation of several state-of-the-art memory systems on HaluMem, including Mem0 (both standard and graph versions) [chhikara2025mem0], Memobase [memobase], MemOS [li2025memos_long], Supermemory [supermemory], and Zep [rasmussen2025zep]. Each memory system was independently evaluated in two subsets, HaluMem-Medium and HaluMem-Long, with efforts made to ensure consistent parameter configurations across evaluations.
@@ 6.2.1 Overall Evaluation on HaluMem
Table 3: Evaluation results of all memory systems on HaluMem. “R” denotes Recall, “Target P” denotes Target Memory Precision, “Acc.” denotes Accuracy, “FMR” denotes False Memory Resistance, “F1” denotes Memory Extraction F1-score, “C” denotes Correct Rate (Accuracy), “H” denotes Hallucination Rate, and “O” denotes Omission Rate. The values in parentheses in the “Target P” and “Acc.” columns represent the number of extracted memories. Color scale reflects performance (red = worse, green = better); Best values in bold.
Dataset System Memory Extraction Memory Updating Question Answering
@@
Medium Mem0 42.91% 65.03% 86.26%(10556) 60.86%(16291) 56.80% 57.31% 25.50% 0.45% 74.02% 53.02% 19.17% 27.81%
Mem0-Graph 43.28% 65.52% 87.20%(10567) 61.86%(16230) 55.70% 57.85% 24.50% 0.26% 75.24% 54.66% 19.28% 26.06%
Memobase 14.55% 25.88% 92.24%(5443) 32.29%(17081) 80.78% 25.13% 5.20% 0.55% 94.25% 35.33% 29.97% 34.71%
MemOS 74.07% 84.81% 86.25%(45190) 59.55%(71793) 44.94% 79.70% 62.11% 0.42% 37.48% 67.23% 15.17% 17.59%
Supermemory 41.53% 64.76% 90.32%(14134) 60.83%(22551) 51.77% 56.90% 16.37% 1.15% 82.47% 54.07% 22.24% 23.69%
Zep - - - - - - 47.28% 0.42% 52.31% 55.47% 21.92% 22.62%
Long Mem0 3.23% 11.89% 88.01%(1134) 46.01%(2433) 87.65% 6.22% 1.45% 0.03% 98.51% 28.11% 17.29% 54.60%
Mem0-Graph 2.24% 10.76% 87.32%(785) 41.26%(1866) 88.36% 4.36% 1.47% 0.04% 98.40% 32.44% 21.82% 45.74%
Memobase 6.18% 14.68% 88.56%(3077) 25.61%(11795) 85.39% 11.55% 4.10% 0.36% 95.38% 33.60% 29.46% 36.96%
MemOS 81.90% 89.56% 82.32%(48246) 43.77%(99462) 28.85% 82.11% 65.25% 0.29% 34.47% 64.44% 16.61% 18.95%
Supermemory 53.02% 70.73% 85.82%(24483) 29.71%(77134) 36.86% 65.54% 17.01% 0.58% 82.42% 53.77% 22.21% 24.02%
Zep - - - - - - 37.35% 0.48% 62.14% 50.19% 22.51% 27.30%
  • Note: since Zep does not provide a Get Dialogue Memory API, metrics related to memory extraction cannot be computed. For details, see Appendix B.
@@
Overall, most memory systems perform worse on HaluMem-Long than on HaluMem-Medium, with Mem0, Mem0-Graph, and Memobase showing particularly notable declines. Notably, the Mem0 series and Memobase extract significantly fewer memories on HaluMem-Long than on HaluMem-Medium, whereas Supermemory and MemOS exhibit the opposite trend. This indicates that future memory systems need to improve their ability to process irrelevant information and distinguish between high- and low-value memories.
In the memory extraction task, regarding memory integrity, except for MemOS, all systems achieve recall (R) rates below 60%, indicating that many reference memory points are not extracted. The higher weighted memory recall (Weighted R) suggests that these systems can prioritize important memory points. Regarding memory accuracy, all systems have accuracy (Acc.) below 62%, reflecting a high proportion of hallucinations, although performance on target memory precision (Target P) is relatively good. Supermemory and MemOS perform the worst on FMR because they tend to extract excessive information without effectively filtering distractions or unhelpful content. Other systems adopt more conservative strategies and thus perform better in FMR. In terms of the F1 score, MemOS and Supermemory perform the best and exhibit stability in long contexts, whereas other systems experience a sharp decline. In summary, future memory systems should strike a balance among coverage of important memories, extraction accuracy, and resistance to interference, aiming for both high quality and reliability in memory retrieval.
In the memory updating task, most systems perform poorly, and their performance drops considerably on HaluMem-Long. Systems showing better performance in memory integrity also tend to exhibit higher update accuracy, but most systems suffer omission rates above 50%. This issue primarily stems from insufficient coverage in memory extraction: when the pre-update memories are not extracted, related updates cannot be properly processed. Moreover, the fact that all systems exhibit hallucination rates below 2% does not necessarily imply strong hallucination suppression, since very few samples actually enter the update stage. Overall, current systems face a clear bottleneck in memory updating: the extraction and updating stages lack stable linkage, resulting in low accuracy and high omission rates.
@@
In the memory question-answering task, the best-performing systems are also those that perform well in memory integrity and memory updating, further highlighting the crucial role of memory extraction. For example, Mem0 and Mem0-Graph show clear performance declines on HaluMem-Long compared to HaluMem-Medium, which strongly correlates with their substantial reduction in extracted memory points. However, all systems achieve answer accuracies below 70%, with both hallucination rate and omission rate remaining high, and their overall performance further decreases on HaluMem-Long. This demonstrates that current memory systems’ QA performance depends heavily on the sufficiency and accuracy of upstream memory extraction, and remains prone to factual deviation and memory confusion under interference or extended context conditions.
6.2.2 Performance on Different Memory Types
Table 4 reports the extraction accuracy of each memory system for event, persona, and relationship memories, which include all memory points from both the memory extraction and updating tasks while excluding distractor memories. MemOS achieves the best overall performance. However, Mem0, Mem0-Graph, and Memobase show a marked decline in long-context scenarios, suggesting difficulty in consistently capturing valuable information in complex dialogues. MemOS and Supermemory perform better on HaluMem-Long than on HaluMem-Medium, probably because they extract a larger number of memory points in the long-context condition. Across memory types, Persona memories yield slightly higher accuracy, indicating that static personal traits are easier to capture, whereas understanding event dynamics and relationship changes remains challenging. Overall, all systems still show low performance across the three memory categories, indicating significant limitations in current memory modeling.
@@ 6.2.3 Performance on Different Question Types
Figure 5 illustrates the performance of different memory systems across six categories of questions. Overall, the accuracy of all memory systems remains relatively low across most categories, indicating substantial room for improvement. The Mem0 series and Memobase show significantly poorer performance on HaluMem-Long compared to HaluMem-Medium, suggesting a notable degradation under ultra-long context conditions. In contrast, MemOS, SuperMemory, and Zep demonstrate relatively stable behavior and achieve consistently superior overall performance on both datasets. Furthermore, all memory systems perform comparatively better on memory boundary and memory conflict questions, indicating their capability to effectively recognize unknown or misleading information and respond correctly. However, their performance deteriorates substantially on multi-hop inference, dynamic update, and generalization & application questions, suggesting that current memory systems still struggle with complex reasoning and preference tracking.
@@ 6.2.4 Efficiency Analysis of Memory Systems
Table 5 shows the time consumption of all memory systems during the evaluation process for dialogue addition and memory retrieval, as well as their total runtime. Overall, dialogue addition requires substantially more time than memory retrieval, indicating that the write stage is the primary computational bottleneck. Enhancing the efficiency of memory extraction and updating is thus crucial for improving interactive performance. On HaluMem-Medium, Supermemory performs best in both dialogue addition and total runtime, while MemOS shows the best retrieval efficiency. However, the dialogue addition time of Mem0 and Mem0-Graph exceeds 2700 minutes, revealing their low processing efficiency during dialogue ingestion and memory construction. On HaluMem-Long, the dialogue addition time for Mem0, Mem0-Graph, and Memobase decreases, mainly because the number of processed memory points is reduced rather than due to performance improvement. In contrast, MemOS and Supermemory extract a substantially larger number of memory points, resulting in a significant increase in their time cost.